To mitigate the issue of ambient light pulluting the QPD signal, I mounted the prototype into a custum built box. This helps a lot. My plan is to add a short piece of black pipe in the front, to further shield from incident light.
The new box also provides a clean way to mount the QPD.
I turned out that all the noise I was seeing in the QPD spectrum was due to ambient light. I covered the QPD with a box and switched off all the light. As shown in the following plot the noise is lower.
Considering that in the final setup we'll have a beam spot radius of 0.5mm, the sensitivity to beam motion on the QPD will be 23e6 V/m. The following plot shows the resulting beam motion sensitivity, if limited by electroninc noise:
It's at a level of 6e-15 m/rHz at all frequencies above 120 Hz.
I got two new ADC and DAC boards from Rolf, with the correct PCIe interface. I installed them into the cymac and checked that the system could boot. The cymac is now sitting in the rack. As requested by Jamie I installed Debian 8.5
Today I gave up trying to fix the first board I populated, and built a second one. The good news is that it's working as expected.
With 27.5 uW incident on each quadrant, I measure about 4.5 V, which is in line with the transimpedance of 200k, a responsivity of about 0.4 A/W and ad additional gain of two coming from the differential driver.
I also measured the noise with a SR785 (it wasn't connected to a GPIB interface and I couldn't find any, so all I have are the following numbers and the attached screenshots).
At low frequency we are dominated by 60 Hz harmonics (probably coming from the laser). At high frequency there are some large peaks of unknown origin. I can't acquire digitally the signals to compute the difference, so I don't know if the noise we see is, for example, laser intensity noise. As soon as the cymac is up and running, I'll run some more tests.
The following table shows the lowest eigenfrequency (Hz) for different sizes of disks
Transimpendance and whitening are working properly. I can't get useful signal out of the differential stages yet. I replaced the channel 1 DRV134 that was burnt (very hot when powered on). But the new one got hot too after powring on, so there might be something else wrong there. I'm also wondering if it's ok to use an oscilloscope to look at the differential stage output. The scope will ground one of the two outputs: according to the DRV134 datasheet this should be ok, but I'll check better later on.
Between yesterday and today I populated one QPD board (based on D1600196), and started testing it. The transimpedance stages seems to work fine (they show about 5-6 V in ambient light). However the whitening stages show a large ~100 kHz oscillation. While trying to fix it I probably burnt one of the output drivers.
I'll continue the investigations and debugging on Monday.
[Massimo Granata (LMA), Quentin Cassar (LMA), Gabriele]
This week I'm visiting LMA to learn how their Gentle Nodal Suspension system works and to measure the quality factors (Q) of one of Mark Optics disks. First of all we annealed the disk for 9 hours at 900 degrees (plus 9 hours warm up and 9 hours cool down).
Then we installed the disk into the measurement system and started by searching for all the resonances.
My COMSOL simulation proved to be good enough to give us the frequencies, especiallty after a small fine tuning of the disk thickness (within specs). We identifies a total of 32 modes of different families, and measured the ring down of all of them. Since our disk has no flats, each mode is actually a doublet with very small frequency separation. The analysis software has a bandwith of 1 Hz to find the peak amplitude, so it can't resolve the two modes. When both are excited to a significant amplitude by the electrostatic actuator, we see a clear beat in the ring-down. I had to write a new fitting code to take this into account. More details will follow in a DCC document. However, here I can say that the fit works remarkably well for all modes.
A couple of examples:
Here is a summary plot of the quality factor and loss angle for all modes. We measured Q as high as 10e6, in line with other LMA samples (2") we tested in these days. In conclusion, the Mark Optics disks, as they are, are good enough for our coating tests.
Instructions on how to setup a workstation are available here:
I'll copy them here and integrate once I got the C.Ri.Me. workstation up and running
** libmotif4 >> libxm4 : sudo apt-get install libxm4
** all .sh files in etc must be modified to point to the correct version of the downloaded software
** add the following line to the end of the ligoapps-userv-end.sh file to get medm and striptool working
** to fix diaggui problem, create a symbolic link in /usr/lib/x86_64-linux-gnu/
sudo ln -s libtiff.so.5 libtiff.so.4
I moved the unused rack from the Crackling Noise lab to the C.Ri.Me lab. It will be used for the new cymac. I also started putting the new workstation together, but I'm missing some adaptors for the monitors.
The PCBs for the QPD circuit and ADC interface are here and look ok. All electronics components are also here (except for the ADC connector which should be ordered separately from Mouser, after confirming that the ADC we're going to use have the same cable as the one we use in the Crackling Noise experiment). The QPD will be shipped on 06/17.
I did some FEA simulation of fused silica disks, to identify the lowest usable eigenmode. By usable I mean a mode that has zero elastic energy stored in the center.
In the attached figures, the dfisk deformation is shown exaggerated, and the color map shows the elastic energy density. All results are obtained with COMSOL/MATLAB, the disk are constrained at a point corresponding to the center of the lower surface. No gravity.
A preliminary design of the ESD board is available on the DCC: D1600214
An improved design is attached. I modified the input telescope to avoid using shor focal length lenses, to make it less critical, and to reduce the beam spot radius at the QPD to 0.5 mm.
Attached a first layout of the optical lever systems. The beam spot radius on the QPD is about 0.8 mm, and the lever arm length is of the orer of 1.4-1.5 m for all four beams.
Here are some screenshots of the disk assembly and a look at how four of them will sit into the vacuum chamber. The Solidworks models are available here: D1600197
The circuit design sent out for fabrication is available in the DCC: D1600196
After a very useful discussion with Rich this morning, I think the circuit based on aLIGO optical levers design should be good for our applications.
It uses a LT1125 as input stage, which has
We expect to send about 5 mW into the disk, getting back a 4% reflection, which would correspond to 200 uW on the QPD. Let's say we lose half of this power in reflections through viewports and such, so we have a total of 100 uW on the QPD, or 25 uW on each quadrant. From the QPD datasheet the repsonse is about 0.4 A/W, so we have a photocurrent of 10 uA. The corresponding shot noise limit is about 1.8e-12 A/rHz.
Using a transimpendace of 200k, the noise at the output of the transimpendance is
So in the worst case the current noise will be about half of the shot noise. This seems good enough.
Koji, Rich, and I recently came up with a new QPD design which is better for general lab use than the aLIGO ones (which have a high-noise preamp copied from iLIGO).
This page has the mechanical drawing only, but perhaps Rich can tell us if he's ready to make the first version for you or not. I think you can get by with the old design, but this new one should be lower noise for low light levels.
For the optical levers we are going to use the same QPD that are used in aLIGO optical levers (see T1600085 and D1100290): Hamamatsu S5981
Based on the aLIGO design, I put together a design fof the QPD boards, see the first attached PDF file. Some comments:
The stable power supply will be provided by an additional board, which will also interface 8 QPD boards to the ADC connector, see the second attached PDF. The ADC signal grounds can be connected directly to the power supply ground, left floating, or connected with a RC filter, depending on what we find to be the best solution.
Total cost estimated for PCB manufacturing and components, including the QPDs is less than 3k$, for a total of 12 boards (we need 8, plus some spares)
We discovered a couple of days ago that the table was sitting on three legs only and the fourth one was dangling. I managed to adjust the height of the fourth leg using the large screw on the leg support. Now the table is properly supported by all four legs.
Elogs for the new Coatin RIng-down MEasurement lab had to start somewhere, so here is a couple of pictures of the optical table with shorter legs and of one of the two vacuum chambers that have been moved in.
Nothing has happened since Steve, the visiting highschool teacher, has left. Meanwhile, some parts of the multi-color BRDF setup were delivered. I assembled everything today and realigned the lasers. Everything is ready now for a three-color BRDF measurement (the previous Richter record was 2 colors). I will claim back my video capture device as soon as possible from my neighbors and then take new images.
We played around with Matlab today. The first step was to convert light wavelengths into RGB colors. In this way we can combine images taken at different colors. The picture shows the purple and red images (stored in gray scale) in heat colormap. Then the sum of these two images is calculated in their natural RGB colors.
Today we improved alignment of the lens-camera arm. We discovered earlier that this alignment affects the amount of "snowfall" on the scattering images. Looking at the latest 405nm video (see attachment), one can still see snowfall, but it is considerably weaker now and the true scatter image is clearly visible. We took a set of scatter images at certain scattering angles and produced BSDF curves. The shape of these curves has partially to do with the snowfall contribution, but one also has to keep in mind that the mirror quality is much worse than what has been used in the Fullerton measurement. We still need to calibrate these curves. The calibration factor is different for the two images so that you cannot even compare them at the moment except for their shape.
Today we also got the new broadband lens for the camera arm. First measurements show that image quality is better. Playing a bit around with distances between object mirror, lens and image plane, we also found that image quality becomes better when the lens and camera get closer to the mirror (which is only an issue for the 405nm measurement since 633nm and 532nm look very good anyway). So we are thinking to change the camera arm setup to make it much shorter.
Here a little purple video. It starts with scattering angle around 15deg and stops at about 80deg.
There are some clear point defects visible especially at small angles.
I will not start to think about some other interesting details of this video before I got the new lens.
Ed: The AVI did not run on Mac. I posted it on youtube. Koji
We have the new 405nm laser pointer. The image to the left shows the scattered light from the red laser, the image to the right scattered light from the purple laser. Both images were taken 30deg with respect to the normal of the mirror surface. Also, we got a new gallon of Methanol. After cleaning the mirror multiple times, the scattered light became significantly weaker. So the purple images look very different from red and green. It could be that the lens that we use to image the mirror surface is the problem since it is specified for the wavelength range 1000nm-1550nm. Could it also be the CCD camera? Anyway, to be sure I will order another broadband lens.
The following two pictures were taken from the same angle with green (left) and red (right) incident laser at an angle of 15deg from the incident beam (reflected to about -5deg). Some scattering centers are collocated. The green laser power is about 5 times as high as the red laser power, but this factor does not seem to calibrate the image well (the green image becomes too dark dividing all pixel values by 5). So there seems to be a significant difference in the divergence of the two lasers. We will have to use a photodiode to get the calibration factor. These images were taken after cleaning the mirror. Before cleaning, there was way too much scattering and the images were mostly saturated.
We were confused a bit about how the camera image changes when you move the arm that holds the camera and lens around the mirror. It seems that scattering centers move in ways that cannot be explained by a misaligned rotation axis. So we wanted to make sure that the mirror surface is actually imaged as we intended to. We generated a white grid with 0.7cm spacing and black background on a monitor. The image that we saw is exactly how we expected it to be. So the image mystery has other reasons.
Steve Maloney, a visiting highschool teacher, and I have started to set up a new scattering experiment in the Richter lab. The idea is to take images of large-angle scattered light using different lasers. We have one 633nm laser, and 532nm and 405nm laser pointers. The goal is to uniformly illuminate the same disk of about 1cm diameter on a silver-coated mirror with all three colors. We use a silver-coated mirror to make sure that the light is reflected from the same layer so that all colors are scattered from the same abberations.
The image shows one of the laser pointers and the HeNe laser. The first step is to widen the beam with a f=5cm broadband, AR coated lens (Newport PAC15AR.15). The diverging beam is then aligned through an iris to give it the right size on the mirror. In this way, illumination is almost uniform on the mirror surface.
The mirror is mounted over the rotation axis of a unipolar stepper motor. For the moment we only took images from fixed direction (initially with a commercial digital camera, later with a monochromatic Sony XT-ST50 CCD camera. The problem with the commercial camera was that you cannot completely control what the camera is doing. Also it would have been very difficult to calibrate the image once you start comparing scattering with different colors. A f=7.5cm lens is used to image the illuminated disk on the CCD chip to make maximal use of its resolution. The CCD signal is read out on a Windows machine with an EasyCap video capture device connected to a USB port. Standard software can then be used to take images or record videos. For some reason the capture device reduces the image size to 640x480 pixels (a little less than the size of the CCD chip).
Eventually the camera and lens will be mounted on a metal arm whose orientation is controlled by the stepper motor. The stepper motor was part of the Silicon Motor Reference Design (Silicon Laboratories). It comes with all kinds of cables and a motor control board. Software is provided to upload compiled C code to the board, but for our purposes it is easiest to use primitive communication methods between the PC and the board. We are working with HyperTerminal that used to be part of Windows installations, but now it has to be downloaded from the web. This program can send simple commands through TCP/IP and COM ports. These commands allow us to position the motor and define its rotation speed. Since our PC does not have a serial port, we purchased a Belkin USB Serial Adapter. You will have to search the web to find suitable drivers for Windows 7 x64. Luckily, Magic Control Technology has similar products and the driver for their U232-P9 USB/serial adapter also works for the Belkin product.
So our goal for the remaining weeks is to take many images from various angles and to set up the experiment in a way that we can VNC into our lab PC and control everything from the Red Door Cafe.